OpenStudy (anonymous):

In a class election with 3 candidates, the winner beat the other 2 candidates by 3 and 6 votes respectively. If 27 votes were cast, how many votes did the winner receive?

OpenStudy (anonymous):

-18 votes

OpenStudy (anonymous):

Call the winner "w". Call the first loser "a" and the second loser "b". w = a + 3 w = b + 6 a + b + w = 27 You now have 3 equations in 3 variables that you should be able to solve. Your question is just looking for "w".

OpenStudy (anonymous):

Hint: Solve the first 2 equations for "a" and "b" respectively and substitute those expressions into the 3rd equation, which will then be one equation in one variable, "w".

OpenStudy (anonymous):

So you mean a+3=b+6 ?

OpenStudy (anonymous):

Not exactly, but your thinking ok about this. You solve the first 2 equations. The first one will be a = something(1) and the second will be b = something(2). You put those "somethings" in to the THIRD equation where you have the "a" and "b".

OpenStudy (anonymous):

I really didn't get what your saying

OpenStudy (anonymous):

Let's look at the first equation. How do you state it in terms of "a" equaling something? Just look at the first equation for now. Re-write it so that the "a" is by itself on one side without the numeric.

OpenStudy (anonymous):

The answer is 12

OpenStudy (anonymous):

Hint: If x + y = z then x = z - y

OpenStudy (anonymous):

You got it!

OpenStudy (anonymous):

Thx a lot. I'll give you a medal for helping me out :)

OpenStudy (anonymous):

Are you from the states?

OpenStudy (anonymous):

I appreciate it! Good luck in your studies! @Chad123

OpenStudy (anonymous):

Yes, right outside Chicago. Shoveled 5 inches of snow today! Gotta love winter! Where you from?

OpenStudy (anonymous):

I m from Melbourne Australia. It's pretty hot down here you would probably want to come over for a holiday to escape the blitz

OpenStudy (anonymous):

I want to get to either Queensland or New Zealand someday. I like the laid-back attitudes I hear about down under.

OpenStudy (anonymous):

You'll love it especially New Zealand and certain parts of Australia. There is a place called Cairns, its always summer up there